a)Round off 76.3 to the nearest integer.
b)Round off 2.9699 to 3 decimal places.
c)Round off 16422 to 3 significant figures.
d)Express 676 as a product of prime factors, giving your answer in index notation.
e)Factorise completely 3a-18b

Answer by richard1234(7193)

a) 76
b) 2.970
c) 16400 (0's are not considered significant figures, unless you have a number such as 5803.7)
d) If you know 676 = 26^2 = (13*2)^2, then 676 factors to 13*13*2*2.
e) 3(a - 6b)